# LUDVIG: Learning-free Uplifting of 2D Visual features to Gaussian Splatting scenes

### v4

@@ 3.3 Enriching features by graph diffusion
DINOv2 features have shown remarkable performance on semantic segmentation tasks with simple linear probing (Oquab et al., 2024), making them a good candidate to enrich features that lack such a property like CLIP (Wysoczańska et al., 2024; Zuo et al., 2024; Liu et al., 2023). Inspired by spectral clustering techniques (Shi & Malik, 2000; Kondor & Lafferty, 2002; Belkin & Niyogi, 2001), we propose to diffuse features that have been uplifted to 3D. This process aims to align semantic features with the scene layout and object boundaries implied by DINOv2. Contrary to the aforementioned methods, we are the first to perform this landscaping with DINOv2 directly in the 3D scene, hereby taking 3D geometry into account as well.
Graph construction
We construct a graph from uplifted DINOv2 features $f$ in $\mathbb{R}^{n}$. Nodes are given by the 3D Gaussians. Edges, represented by a matrix $A$ of size $n\times n$, encode both the 3D Euclidean geometry between the nodes and the similarity between their DINOv2 features. More precisely, we extract the $k$ nearest neighbors $\mathcal{N}(i)$ for each node $i$, as measured by the Euclidean distance between the centers of the 3D Gaussians. Two nodes $i$ and $j$ in the graph are linked by an edge if $i\in\mathcal{N}(j)$ or $j\in\mathcal{N}(i)$, and the edge is assigned the following weight:
$A_{ij}=S_{f}(f_{i},f_{j})\,P(f_{i})^{\frac{1}{2}}P(f_{j})^{\frac{1}{2}},$ (7)
with $S_{f}(f_{i},f_{j})$ a local similarity between features $f_{i}$ and $f_{j}$, typically defined as a RBF kernel. For tasks requiring diffusion to be confined to a specific object instance, we prevent leakage into the background by introducing a node-wise unary regularization term $P(f_{i})$ which quantifies the similarity between the node feature $f_{i}$ and the features of the object of interest. Details on $S_{f}$ and $P$ are provided in Appendix  A.3.
Diffusion on the graph.
Given initial 3D features $g_{0}$ in $\mathbb{R}^{n}$, which we aim to improve by using information encoded in $A$ (3D geometry and DINOv2 similarities), we perform $T$ diffusion steps to construct a sequence of diffused features $(g_{t})_{1\leq t\leq T}$ defined as follows:
$g_{t+1}=A\tilde{g}_{t},\quad\tilde{g}_{t}={g_{t}}/{\|g_{t}\|_{2}},$ (8)
This can be seen as performing a few steps of the power method, making $g_{0}$ closer to the dominant eigenspace of $A$. Note that depending on the downstream task, $g_{0}$ may represent generic features or task-specific features such as coarse 3D segmentation masks.
@@ 4 From 3D Uplifting to downstream tasks
In this section, we describe our approach for extracting and uplifting features from SAM, DINOv2, and CLIP, as well as evaluating these features on two downstream tasks: segmentation and open-vocabulary object localization. As in Sec. 3, we are given a set of 2D frames $I_{1},\dots,I_{m}$, with viewing directions $d_{1},\dots,d_{m}$, and a 3D Gaussian Splatting representation of the scene.
@@ 4.1 Multi-view Segmentation
We assume that a foreground mask of the object to be segmented is provided on the reference frame $I_{1}$. The foreground masks are either scribbles or a whole reference mask of the object, both of which define a set of foreground pixels $\mathcal{P}$. In the following, we present the proposed approaches for segmentation using SAM and DINOv2 features, based on both types of foreground masks.
Multi-view segmentation with SAM.
SAM (Kirillov et al., 2023; Ravi et al., 2024) is a powerful model that can generate object segmentation masks from point prompts, on a single 2D image. Aggregating SAM 2D segmentation masks in 3D allows for cross-view consistency and improves single-view segmentation results. We proceed by generating 2D feature maps based on SAM segmentation masks of each training frame while only relying on the foreground mask for the reference frame $I_{1}$. The 2D feature maps are generated by constructing several sets of point prompts on each training frame which are then provided to SAM to obtain several segmentation masks. The point prompts are obtained using the foreground mask provided on the reference frame as described in Appendix A.1. Averaging the resulting segementation masks for each frame results in the final 2D SAM feature maps. These are then uplifted using the aggregation scheme in Sec. 3.2. Our final prediction is obtained by rendering the uplifted feature maps into the target frame and thresholding.
Multi-view segmentation with DINOv2.
We construct 2D feature maps at the patch level using DINOv2 with registers (Darcet et al., 2024) and uplift them into a high resolution and fine-grained 3D semantic representation which is then used for segmentation. The 2D feature maps are constructed using a combination of a sliding windows mechanism and dimensionality reduction of the original DINOv2 features as described in Appendix A.2 and illustrated in Figure 4 therein. This approach enhances the granularity of spatial representations by aggregating patch-level representations to form pixel-level features. To favor the first principal components, known to focus on the foreground objects (Oquab et al., 2024), the features are re-weighted by the eigenvalues of the PCA decomposition. The 2D feature maps from the $m$ training views are uplifted using Eq. (3) and the resulting 3D features are then re-projected into any viewing direction $d$ using Eq. (4) to compute rendered 2D features $(\hat{F}_{d,p})$. To obtain segmentation masks, we define a predictor score $P(\hat{F}_{d,p})$ as the likelihood that a 2D pixel $p$ belongs to the foreground, based on its feature $\hat{F}_{d,p}$. The score $P$ is obtained by comparing the rendered features $(\hat{F}_{d,p})$ with foreground features $\mathcal{F}_{\text{ref}}:=(\hat{F}_{d_{1},p})_{p\in\mathcal{P}}$ corresponding to the foreground mask from the reference frame $I_{1}$, and the final segmentation mask is then obtained by thresholding. More details are provided in Appendix A.2.
Enhancing segmentation with DINOv2 using 3D graph diffusion.
DINOv2 provides generic visual features that do not explicitly include segmentation information, unlike models such as SAM that were specifically trained for such a task. Consequently, 2D projections of uplifted DINOv2 features might fail to separate distinct objects that have similar features. This challenge can be mitigated by incorporating 3D spatial information.
@@
To this end, we propose to leverage the graph diffusion process introduced in Section 3.3 and illustrated in Figure 3. We set the initial vector of weights $g_{0}\in\mathbb{R}^{n}$ of the graph diffusion algorithm to be a coarse estimation of the contribution of each Gaussian to the final segmentation mask. This initial weight vector is computed by uplifting the 2D foreground mask (either scribbles or a reference mask) from the reference frame into 3D using Eq. (3), normalizing and thresholding them (see Appendix A.3) The nodes for which $g_{0}$ has a positive value define a set of anchor nodes $\mathcal{M}$ that are more likely to contribute to the foreground. The regularization term $P$ appearing in Eq. (7) is obtained by comparing the uplifted features with anchor features obtained using the foreground mask as described in the appendix. For this task, we binarize $A$ with a fixed threshold (set to $10^{-5}$). After the $T$ diffusion steps, we recover the nodes $\mathcal{S}$ in $g_{T}$ with strictly positive values (i.e., those reachable after $T$ iterations). The final weight is defined as $h_{i}=P(f_{i})$ if $i\in\mathcal{S}$ and 00 otherwise. Segmentation is then performed by projecting $\mathbf{h}=(h_{i})$ into 2D and thresholding.
4.2 Open-vocabulary object detection
@@
Following Kerr et al. (2023); Qin et al. (2024); Zuo et al. (2024), we uplift CLIP features (Ilharco et al., 2021) for open-vocabulary object localization. CLIP effectively aligns images and text in a shared representation space. As a measure of alignment, we use the relevancy score introduced by LERF (Kerr et al., 2023), which quantifies the similarity between a CLIP visual feature and a text query.
Construction of CLIP feature maps.
@@
We follow common practice (Kerr et al., 2023; Zuo et al., 2024) and construct multi-resolution CLIP 2D feature maps by querying CLIP on a grid of overlapping patches at different scales and aggregating the resulting representations. As in Zuo et al. (2024), rather than keeping the different representations separate, we aggregate them with a simple average pooling. These multi-resolution CLIP features are uplifted into 3D using Eq. (3).
Relevancy scores.
@@
After uplifting CLIP features, we compute relevancy scores for each Gaussian’s feature to text queries embedded by CLIP. These relevancy scores can then be projected into 2D and used for both localization and segmentation. For localization, we choose the pixel with the highest relevancy score. For segmentation, we predict a SAM mask by selecting point prompts among pixels with the highest relevancy score. Specifics on the computation of relevancy scores and segmentation masks are provided in Appendix A.4.
Refining relevancy scores with DINOv2 graph diffusion.
@@
We refine 3D relevancy scores with the diffusion process described in Sec. 3.3. To this end, DINOv2 features are also uplifted, and the similarity matrix is built as in Eq. (7), with the unary term $P$ constructed using a logistic regression over thresholded relevancies, see details in Appendix A.4. The diffusion process propagates CLIP relevancies to Gaussians with similar DINOv2 features. The resulting 3D relevancy scores span the object of interest without covering other objects with similar features and show a strong decay at the object’s borders, as defined by DINOv2’s feature landscape, resulting in improved segmentation results.
@@ 5 Experiments
5.1 Experiment details
3D scene training and pruning.
All scenes are trained using the original Gaussian Spatting implementation (Kerbl et al., 2023) with default hyperparameters. For memory efficiency, half of the Gaussians are filtered out based on their importance, as described in Sec. 3.2.
2D vision models.
Our experiments are conducted using DINOv2’s ViT-g with registers (Darcet et al., 2024), SAM (Kirillov et al., 2023), SAM 2 (Ravi et al., 2024) and the OpenCLIP ViT-B/16 model (Ilharco et al., 2021).
Segmentation tasks.
We consider two segmentation tasks: i) Neural Volumetric Object Selection (NVOS,  Ren et al. 2022), which is derived from the LLFF dataset (Mildenhall et al., 2019), and ii) SPIn-NeRF, which contains a subsets of scenes from NeRF-related datasets (Knapitsch et al., 2017; Mildenhall et al., 2019; 2021; Yen-Chen et al., 2022; Fridovich-Keil et al., 2022). The NVOS dataset consists of forward-facing sequences where the task is to predict the segmentation mask of a labeled frame based on reference scribbles from another frame. SPIn-NeRF contains both forward-facing and 360-degree scenes, in which all frames are labeled with segmentation masks, and the standard evaluation protocol uses the segmentation mask from the first frame as reference to label the subsequent frames.
Open-vocabulary object localization
We evaluate our approach on the LERF dataset (Kerr et al., 2023), which includes localization and segmentation tasks on complex in-the-wild scenes. We report our results on the extended evaluation task introduced by LangSplat (Qin et al., 2024) containing additional challenging localization samples, and also follow their evaluation protocol.
Evaluation and hyperparameter tuning.
Results are averaged over 3 independent runs, and hyperparameters are selected as follows. Segmentation with 3D SAM masks requires setting a threshold for foreground/background pixel assignment, and optionally choosing one of the three masks proposed by SAM (representing different possible segmentations of the object of interest). Segmentation with DINOv2 relies on two parameters for defining $S_{f}$ and $P$ in graph diffusion, and on the threshold for foreground/background pixel assignment. These parameters are automatically chosen as described in Appendix Secs. A.1 and A.3.
For the open-vocabulary object localization tasks, all parameters including those used for graph diffusion and thresholds for segmentation are either fixed for all scenes or automatically selected, as detailed in Appendix A.4.
@@ 5.1 Experiment details
(b) Single-view DINOv2 PCA
(c) Multi-view DINOv2 PCA
@@
Figure 3: Illustration of the diffusion process. 2D projection of the weight vector $g_{t}$ (white) and unary regularization term (red) at different diffusion steps $t$. The diffusion process allows filtering out unwanted objects that have similar features to the object of interest (such as the two smaller skulls on horns, bottom-row), but are disconnected in space. The regularization term (red background) prevents leakage from the object to the rest of the scene (such as through the fern’s trunk, top-row).
5.2 Qualitative results
DINOv2 feature uplifting.
First, we illustrate the effectiveness of our simple uplifting approach. Figure 2 shows the first three PCA components (one channel per component) over DINOv2’s patch embeddings. The coarse patch-level representations from every view (middle) are aggregated using Eq. (5) to form a highly detailed 3D semantic representation, and reprojected into 2D (right) using Eq. (4). The aggregation is very fast, being directly implemented in the Gaussian Splatting CUDA-based rendering process, and takes about 1.3ms per view and feature dimension. The first principal component (encoded in the red channel) mostly captures the foreground object, and the subsequent ones allow refining the foreground representations and delivering a detailed background. In the appendix, we provide additional comparative visualizations of our learned 3D features (Figure 8) and of 3D segmentation for scene editing (Figure 7).
Graph diffusion.
Figure 3 illustrates the effectiveness of the diffusion process. The graph nodes are initialized with the reference scribbles, and the diffusion spreads through the object of interest, stopping at its borders. As illustrated with the case of Horns, diffusion filters out unwanted objects that are similar to the object of interest (here, the two skulls on the side). In the Fern scene, diffusion progressively spreads through the branches to their extremities, with the regularization (red background) constraining it to the trunk and preventing leakage, even after a large number of iterations. Appendix Figure 6 also illustrates this for the Flower and Trex scenes: diffusion rapidly spreads, achieving near-full coverage after only 5 steps before reaching all the much smaller Gaussians on the border, allowing for a refined segmentation.
NVOS MVSeg SA3D-TRF SA3D-GS SAGA OmniSeg3D LUDVIG (Ours)
3D representation: TensoRF GS GS NeRF GS
Uplifting: SAM SAM SAM SAM DINOv2 SAM SAM2

### v5

@@ 3.3 Enriching features by graph diffusion
DINOv2 features have shown remarkable performance on semantic segmentation with simple linear probing [33], making them a good candidate to enrich features that lack such a property, such as those from CLIP [47, 56, 28]. Inspired by spectral clustering techniques [40, 24, 1] and manifold denoising [14], we propose to diffuse features that have been uplifted to 3D. This process aims to align semantic features with the scene layout and object boundaries implied by DINOv2. In contrast to prior work, we perform this landscaping with DINOv2 directly in the 3D scene, thereby taking 3D geometry into account as well.
Graph construction. We construct a graph whose nodes are the $n$ 3D Gaussians and edges, represented by a matrix $A$ of size $n\times n$, are based on 3D Euclidean geometry between the nodes and the similarity between their DINOv2 features. More precisely, we extract the $k$ nearest neighbors $\mathcal{N}(i)$ for each node $i$, as measured by the Euclidean distance between the centers of the 3D Gaussians. Two nodes $i$ and $j$ in the graph are linked by an edge if $i\in\mathcal{N}(j)$ or $j\in\mathcal{N}(i)$, and the edge is assigned the following weight:
$A_{ij}=S_{f}(f_{i},f_{j})\,P(f_{i})^{\frac{1}{2}}P(f_{j})^{\frac{1}{2}},\vskip-1.9919pt$ (7)
with $S_{f}(f_{i},f_{j})$ a local similarity between features $f_{i}$ and $f_{j}$, typically defined as a RBF kernel. For tasks requiring diffusion to be confined to a specific object instance, we prevent leakage into the background by introducing a node-wise unary regularization term $P(f_{i})$ which quantifies the similarity between the node feature $f_{i}$ and the features of the object of interest. Details on $S_{f}$ and $P$ are provided in the supplementary (Sec. A.3).
Diffusion on the graph. Given initial 3D features $g_{0}$ in $\mathbb{R}^{n}$, which we aim to improve by using information encoded in $A$ (3D geometry and DINOv2 similarities), we perform $T$ diffusion steps to construct a sequence of diffused features $(g_{t})_{1\leq t\leq T}$ defined as follows:
$g_{t+1}=A\tilde{g}_{t},\quad\tilde{g}_{t}={g_{t}}/{\|g_{t}\|_{2}},\vskip-1.9919pt$ (8)
This can be seen as performing a few steps of the power method, projecting $g_{0}$ into the dominant eigenspace of $A$. Depending on the task, $g_{0}$ may represent generic features or task-specific features such as 3D segmentation masks.
@@ 4 From 3D uplifting to downstream tasks
In this section, we describe our pipeline around the uplifting process from Sec. 3, covering preceding and following stages for three tasks: multi-view segmentation, open-vocabulary object segmentation, and open-vocabulary semantic segmentation. As in Sec. 3, we are given a set of 2D frames $I_{1},\dots,I_{m}$, with viewing directions $d_{1},\dots,d_{m}$, and a 3D Gaussian Splatting representation of the scene.
@@ 4.1 Multi-view segmentation
We assume that a foreground mask of the object to be segmented is provided on the reference frame $I_{1}$. This mask can either be a full segmentation or a sparse set of scribbles, both defining a set of foreground pixels $\mathcal{P}$. The task is to generate a 2D segmentation mask on one or more target frames based on the foreground from the reference frame. It is evaluated with the intersection over union (IoU).
Segmentation with SAM. SAM [21, 37] is a supervised pretrained model that, given an input image and point prompts, generates segmentation masks. We uplift SAM masks from multiple views using Eq. (3); this aggregation improves cross-view consistency and leads to better single-view segmentation. On each view, SAM is run multiple times with different prompt sets obtained by uplifting and reprojecting the reference mask, and the resulting predictions are averaged (see Sec. A.1). The final prediction is obtained by rendering the 3D mask into the target frame and thresholding. Results obtained with this strategy are reported for SAM [21] and SAM2 [37] in Sec. 5.2.
Segmentation with DINOv2. We construct 2D feature maps using DINOv2 with registers [9], applying a sliding window over the image followed by dimensionality reduction of the DINOv2 features. To favor the first principal components, known to focus on the foreground objects [33], the features are re-weighted by the eigenvalues of the PCA decomposition. The 2D feature maps from the $m$ training views are uplifted using Eq. (3) and the resulting 3D features are re-projected into any direction using Eq. (4). 2D segmentation is obtained by thresholding the similarity between projected features and those from the foreground pixels $\mathcal{P}$ in the reference frame (see Sec. A.2). This approach corresponds to an ablation discussed in Sec. 5.2.
@@
Improving DINOv2 segmentation with graph diffusion. Segmentation based on DINOv2 alone may struggle to distinguish objects with similar features, as it lacks 3D spatial information. To address this, we leverage the graph diffusion process introduced in Sec. 3.3 and illustrated in Fig. 4. The foreground mask from the reference frame is first uplifted into 3D, defining a set of anchor Gaussians $\mathcal{M}$. This set is used to initialize the weight vector $g_{0}$ and to define the regularization term $P$, based on the similarity between each 3D feature and those of Gaussians in $\mathcal{M}$ (see supplementary Sec. A.3). For this task, we binarize $A$ with a fixed threshold (set to $10^{-5}$). After $T$ diffusion steps, we recover the nodes $\mathcal{S}$ in $g_{T}$ with strictly positive values (i.e., those reachable after $T$ steps). The final weight is defined as $h_{i}=P(f_{i})$ if $i\in\mathcal{S}$ and 0 otherwise. Segmentation is then performed by projecting $\mathbf{h}=(h_{i})$ into 2D and thresholding. This approach achieves results comparable to SAM, as reported and discussed in Sec. 5.2.
4.2 Open-vocabulary object segmentation
@@
Following [19], we tackle the task of open-vocabulary object localization by uplifting CLIP features [15]. CLIP effectively aligns images and text in a shared representation space. As a measure of alignment, we use the relevancy score introduced by LERF [19], which measures the similarity between a CLIP visual feature and a text query.
@@
Construction of CLIP feature maps. We follow common practice [19, 56] and construct multi-resolution CLIP 2D feature maps by querying CLIP on a grid of overlapping patches at different scales. As in [56], rather than keeping the resulting representations separate, we aggregate them via average pooling. These multi-resolution CLIP features are uplifted into 3D using Eq. (3).
@@
Relevancy scores. After uplifting CLIP features, we compute relevancy scores for each Gaussian’s feature to text queries embedded by CLIP. These relevancy scores can then be projected into 2D and used for both localization and segmentation. For localization, we choose the pixel with the highest relevancy score. For segmentation, we render the relevancy scores and either (i) directly threshold the rendered scores, or (ii) predict a SAM mask by selecting point prompts among pixels with the highest score. Specifics on the computation of relevancy scores and segmentation masks are provided in the supplementary (Sec. A.4).
@@
Refining relevancy scores with DINOv2 graph diffusion. We refine 3D relevancy scores with the diffusion process described in Sec. 3.3. To this end, DINOv2 features are also uplifted, and the similarity matrix is built as in Eq. (7), with the unary term $P$ constructed using a logistic regression over thresholded relevancies, see details in the supplementary (Sec. A.4). The diffusion process propagates CLIP relevancies to Gaussians with similar DINOv2 features. The resulting 3D relevancy scores span the object of interest without covering other objects with similar features and show a strong decay at the object’s borders, as defined by DINOv2’s feature landscape, resulting in improved segmentation results. This approach is evaluated and compared to direct segmentation without diffusion in Sec. 5.3.
4.3 Open-vocabulary semantic segmentation
In semantic segmentation, one needs to label each Gaussian with the appropriate text label from a predefined set. For this task, we directly replace OpenGaussian’s 3D feature learning with our uplifting approach, keeping the rest of their protocol unchanged. This enables a fair comparison under the same evaluation setup. In particular, 2D feature map generation is performed using SAM in automatic mask generation mode and assigning a CLIP feature to each mask, as in [35, 41, 46]. After uplifting, each Gaussian is assigned the text label with the highest CLIP similarity. Results are reported in Sec. 5.4.
@@ 5 Experiments
In the following experiments, we train all scenes using the original Gaussian Splatting implementation [18] with default hyperparameters. To reduce memory usage, half of the Gaussians are filtered out based on their importance, as described in Sec. 3.2.
We uplift features from DINOv2 ViT-g with registers [9], SAM [21], SAM 2 [37], and OpenCLIP ViT-B/16 [15]. Key parameters (e.g. the segmentation threshold and the definitions of $S_{f}$ and $P$ in graph diffusion) are either fixed across all scenes of a task or automatically selected, as detailed in the supplementary material (Secs. A.1, A.3, A.4). Results are averaged over three independent runs.
5.1 Qualitative results
DINOv2 feature uplifting.
First, we illustrate the effectiveness of our simple uplifting approach. Fig. 3 shows the first three PCA components (one channel per component) over DINOv2’s patch embeddings. The coarse patch-level representations from every view (middle) are aggregated using Eq. (5) to form a highly detailed 3D semantic representation, and reprojected into 2D (right) using Eq. (4). The aggregation is very fast, being directly implemented in the Gaussian Splatting CUDA-based rendering process, and takes about 2ms per view and feature dimension. The first principal component (encoded in the red channel) mostly captures the foreground object, and the subsequent ones allow refining the foreground representations and delivering a detailed background. In the supplementary, we provide additional comparative visualizations of our learned 3D features and of 3D segmentation for object removal.
Graph diffusion. Fig. 4 illustrates the diffusion process. The graph nodes are initialized with the reference scribbles, and the diffusion spreads through the object of interest. As illustrated with the case of Horns, diffusion filters out unwanted objects that are similar to the object of interest (here, the two skulls on the side). In the Fern scene, diffusion progressively spreads through the branches to their extremities, with the regularization (red background) constraining it to the trunk and preventing leakage, even after a large number of iterations. Fig. C in the supplementary also illustrates this for the Flower and Trex scenes: diffusion rapidly spreads, achieving near-full coverage after only 5 steps before reaching all the much smaller Gaussians on the border, allowing for a refined segmentation.
Open-vocabulary object removal. Fig. 1 and 5 provide qualitative results for the object removal task, performed by discarding all Gaussians corresponding to a binary 3D segmentation mask. This mask is obtained by computing relevancy between 3D CLIP features and the scene-specific textual prompt “bonsai in a ceramic pot” or “teddy bear”, followed by DINOv2-guided graph diffusion.
@@ 5.1 Qualitative results
(b) Single-view PCA
(c) Multi-view PCA
@@
Figure 4: Diffusion process. 2D projection of the weight vector $g_{t}$ (white) and unary regularization term (red) at diffusion steps $t$. The diffusion process filters out unwanted objects that have similar features to the object of interest (such as the two smaller skulls on horns, bottom-row), but are disconnected in space. The regularization term (red background) prevents leakage from the object to the rest of the scene (such as through the fern’s trunk, top-row).
Figure 5: Open-vocabulary object removal. Removing the teddy bear from the CO3D dataset [38], using DINOv2-guided graph diffusion based on 3D CLIP relevancy scores.
5.2 Multi-view segmentation results
Task. In this section, we consider two segmentation tasks: (i) Neural Volumetric Object Selection (NVOS) [39], which is derived from the LLFF dataset [30], and (ii) SPIn-NeRF, which contains a subsets of scenes from NeRF-related datasets [22, 30, 31, 51, 11]. NVOS consists of forward-facing sequences where the task is to predict the segmentation mask of a labeled frame based on reference scribbles from another frame. SPIn-NeRF contains both forward-facing and 360-degree scenes, in which all frames are labeled with segmentation masks, and the standard evaluation protocol uses the segmentation mask from the first frame as reference to label the subsequent frames.
Setup. We evaluate segmentation based on SAM and SAM2 mask uplifting and on DINOv2 feature uplifting combined with graph diffusion (see Sec. 4.1). We compare our segmentation results (IoU) to the current state of the art: SA3D [4], SA3D-GS [3], SAGA [5], OmniSeg3D [52]. All these methods are specifically designed for uplifting the 2D segmentation masks produced by SAM into 3D using gradient-based optimization of a projection loss. We also report results from NVOS [39] and MVSeg [51], who respectively introduced the NVOS and SPIn-NeRF datasets.
Results. Table 1 reports the average IoU across all scenes for NVOS and SPIn-NeRF. Per-scene results can be found in supplementary Tables C and D. Our results are comparable to the state of the art, while not relying on gradient-based optimization. Surprisingly, our segmentation with DINOv2 using graph diffusion also gives results on par with models leveraging SAM masks. Compared to SAM, DINOv2 better captures complex objects but sometimes also captures background noise. This can be seen in supplementary Fig. B, T-Rex example: while SAM misses the end of the tail and ribs, DINOv2 captures the whole T-Rex but also includes part of the stairs behind. Our lower segmentation results compared to OmniSeg’s can partly be attributed to the poor Gaussian Splatting reconstruction of highly specular scenes, such as Fork. As also noted by [5], the reconstruction includes semi-transparent Gaussians floating over the object, attempting to represent reflections or surface effects, which are challenging to capture using standard rasterization techniques [16].
Ablation study. We compare our segmentation protocol using DINOv2 and SAM2 to multiple simpler variants. More precisely, we evaluate (i) a purely geometrical variant that reprojects the reference mask on the other views, without using SAM2 or DINOv2, (ii) single-view segmentation in 2D based on SAM2 or DINOv2 2D predictions, (iii) uplifting DINOv2 features or SAM2 masks into 3D then rendering them for segmentation, and (iv) segmenting using graph diffusion over DINOv2 3D feature similarities. Results are reported in Table 2, and per-scene IoU as well as a detailed analysis can be found in supplementary Table E and Sec. C.1. We observe that the purely geometrical approach works well on the forward-facing scenes and fails on 360-degree scenes. The single-view variant performs reasonably well on average, but the low resolution of patch-level representations (illustrated in Fig. 3) lead to a coarser segmentation. 3D uplifting considerably boosts results compared to single-view approaches, and introducing 3D spatial information through 3D graph diffusion further enhances results on the more challenging 360-degree scenes.
NVOS [39] MVSeg [32] SA3D-TRF [4] SA3D-GS [3] SAGA [5] OmniSeg3D [52] LUDVIG (Ours)
3D representation TensoRF 3D-GS 3D-GS NeRF 3D-GS
Uplifting SAM SAM SAM SAM DINOv2 SAM SAM2